An air-filled parallel-plate capacitor is to be constructed which can store 12 μC of charge when operated at 1200 V. What can be the minimum plate area of the capacitor? The dielectric strength of air is 3 × 10^6 V/m.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. Q = 12 μc V = 1200 V v/d = 3 * ^10-6 v/m d = v/(v/d) = 1200/3 * 10^-6 = 4 * 10^-4 m c = Q/v = 12 * 10^-6/1200 = 10^-8 f ∴ C = ε base 0A/d = 10^-8 f ⇒ A = 10^-8 * d/ε base 0 = 10^-8 *4 * 10^-4/8.854 * 10^-4 0.45m^2